find the x and y intercepts:
y = x squared + 8x + 14
: find the x and y intercepts:
y = x squared + 8x + 14

Answer by Mathtut(1352)

this is a parabola the y intercept is found setting x's value to zero
so y intercept is 14 or (0,14)
the x intercepts are found by setting y=0
so we have x^2+8x+14 using the quadratic formula we come up with x intercepts of -2.59 and -5.41 or (-2.59,0)and (-5.41,0)
Quadratic equation (in our case ) has the following solutons:


For these solutions to exist, the discriminant should not be a negative number.
First, we need to compute the discriminant : .
Discriminant d=8 is greater than zero. That means that there are two solutions:




Quadratic expression can be factored:
Again, the answer is: -2.58578643762691, -5.41421356237309. :

Question 167601: find the x and y intercepts:
y = x squared + 8x + 14
: find the x and y intercepts:
y = x squared + 8x + 14

Answer by Alan3354(1931)

find the x and y intercepts:
y = x squared + 8x + 14
-----------------
To find the y-intercept, set x=0
y = 14, so it's (0,14)
Finding the x-intercepts is solving for x
Solved by pluggable solver: SOLVE quadratic equation (work shown, graph etc)
Quadratic equation ax^2+bx+c=0 (in our case 1x^2+8x+14 = 0) has the following solutons:

x[12] = (b+-sqrt( b^2-4ac ))/2\a

For these solutions to exist, the discriminant b^2-4ac should not be a negative number.

First, we need to compute the discriminant b^2-4ac: b^2-4ac=(8)^2-4*1*14=8.

Discriminant d=8 is greater than zero. That means that there are two solutions:  x[12] = (-8+-sqrt( 8 ))/2\a.

x[1] = (-(8)+sqrt( 8 ))/2\1 = -2.58578643762691
x[2] = (-(8)-sqrt( 8 ))/2\1 = -5.41421356237309

Quadratic expression 1x^2+8x+14 can be factored:
1x^2+8x+14 = (x--2.58578643762691)*(x--5.41421356237309)
Again, the answer is: -2.58578643762691, -5.41421356237309. Here's your graph:
graph( 500, 500, -10, 10, -20, 20, 1*x^2+8*x+14 )

To clarify, the x-intercepts are:
-4 + sqrt(2) and
-4 - sqrt(2)